How many three digit numbers greater than 700 can be formed by using the digits 4,5,6,7,8,9?

Answer by mathstutor458(57)

given numbers are 4,5,6,7,8,9
the required number is greater than 700 and that must be a three digit number.
first number must be 7,8,9
the remaining numbers can be any remaining numbers.
Therefore,numbers greater than 700 =3*6*6=108(if numbers are repeting,if numbers are not repeting then answer=3*5*4=60)